The month rent of a flat this year is $32000 less than three times the monthly rent of last year. If the monthly rent this year is $4000 more than that of last year, find the monthly rent of last year and this year.
Answer by ikleyn(52781)

.

Let x = month rent this years;

    y =   that  of last year.


Then from condition, you have these equations

    x = 3y - 32000    (1)

    x = y  + 4000     (2)


The rest is just routine exercise.


For example, from the equations you immediately get


    3y - 32000 = y + 4000,


and then


    2y = 32000 + 4000 = 36000

     y                = 36000/2 = 18000


and so on . . . 

----------

The numbers are not realistic, but I think that the goal is more to teach/(to learn) making setup than getting (realistic) solution.

If so, then the author/(the creator) deserves to be forgiven.